2x+1=51 is one of the equasions on my nephews question sheet it has been 30 years since i was in school and i cant exsplain the answer to him please help me remember

Answer by Flannery(124)

2x+1=51
2x=50 (subtract one from both sides)
x=25 (divide both sides by two)
remember in each step to do the some operation on both sides of the equation
keep plugging till you isolate x
here's the check for the problem
2(25)+1=51
50+1=51
51=51
Q.E.D.